OpenStudy (anonymous):

let b represent Barry's age and c Clair's age. Clair is 6 years older than Barry. Solve and express the following in the terms of b and or c. 1. Barry's age in 4 years 2. Clair's age 2 years ago 3. The sum of their ages 4 years ago 4. The sum of their ages 4 years hence 5. Barry is half as old as Clair 6. Clair is thrice as old as Barry 7. Four years ago, Barry was one-third as old as Clair was 8. Next year, the sum of Barry's age and Clair's age will be 26

OpenStudy (anonymous):

1. 4 2. c-2 3. b-4+c-4 4. b+4+c+4=b+c+8 5. b=(c/2) 6. 3b=c 7. 3(b-4)=c-4 OR 3b - 12 = c -4 OR 3b - c - 8 = 0 8. b+1 + c + 1 = 26 OR b + c = 24

OpenStudy (anonymous):

tnx ..:)) what is the value of b and c? thank you so much..

OpenStudy (anonymous):

changes every time. you can only ever express one's age as a function of the other's